sinx *siny=1\4 3tanx=4tany prove that sin(x+y)=7\16

sinx *siny=1\4
3tanx=4tany
prove that
sin(x+y)=7\16

Grade:12

2 Answers

Arun
6 years ago
Dear manish
 
your question is wrong. Please check and repost it or you can also attach an image.
I think it hsould be sin x cos y = 1/4
 
regards
Arun (askIITians forum expert)
rajeswari
6 years ago
given that 3/4=tanx/tany
3/4=(sinx*cosy)/(cosx.siny)
(cosx*siny)*3/4=sinx*cosy
sin(x+y)=sinxcosy+cosxsiny
sin(x+y)=sinxcosy(1+3/4)
sin(x+y)=(1/4)*(7/4)
sin(x+y)=7/16